# FuxiTranyu: A Multilingual Large Language Model Trained with Balanced Data

### v2

@@ 3.4 Model Architecture
  • Normalization. Given the significance of pre-training stability in training large LMs with a substantial number of tokens, we implement pre-normalization due to its superior stability compared to post-normalization Xiong et al. (2020). Furthermore, we incorporate the widely used RMSNorm Jiang et al. (2023b) to enhance training efficiency.
@@
# Params 8B
Hidden Size 4,096
Intermediate Size 16,384
Heads 32
Layers 30
Position Embed 4,096
Vocab Size 250,752
Learning Rate 3e-4 $\rightarrow$ 1e-4
Batch Size 2M $\rightarrow$ 4M
Context Length 4,096
Training Tokens 606B
FlashAttn V2 $\checkmark$
3.5 Pre-training Details
The training procedure for the FuxiTranyu model adheres to the standard autoregressive language model framework, utilizing the next-token prediction loss as detailed in (Brown et al., 2020). To enhance pre-training efficiency, we employ a document packing method similar to that described in (Raffel et al., 2020). This involves randomly shuffling documents, merging them, and then truncating into multilingual chunks that adhere to a maximum context length of 4096 tokens during the pre-training phase.
To mitigate memory consumption and further improve training efficiency, we leverage ZeRO-2 Rajbhandari et al. (2020) and Flash-Attention V2 Dao (2024) technologies. For optimization, the standard AdamW optimizer Loshchilov and Hutter (2017) is utilized with hyper-parameters set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$. We employ the cosine learning rate scheduler, starting with a maximum learning rate of 3e-4 and decaying to a minimum of 10% of the maximum rate. Notably, after encountering divergence issues post-training approximately 241 billion tokens, we reduced the maximum learning rate to 1e-4 to match with the learning rate used in BLOOM, given the multilingual context of both models.
@@ 4 Post-training
To develop a model capable of following instructions and engaging in conversational interactions with humans, we have adopted the instruction fine-tuning and reinforcement learning (RL) approach outlined in (Ouyang et al., 2022).
During the instruction fine-tuning phase, we curate a diverse and high-quality open-source instruction dataset. Given the abundance of instruction-following datasets that have demonstrated exceptional alignment results with various models, manually selecting and fine-tuning the mixture rates for each dataset becomes a challenging task. Consequently, we opt to designate a primary dataset and supplement it with additional datasets. In this context, we select the OpenHermes 2.5 data collection Teknium (2023) as our base dataset, composed of multiple datasets covering a wide range of instructions and yielding excellent results when fine-tuned with Mistral-7B-v0.1. We make modifications to the original OpenHermes 2.5 dataset by replacing Airoboros 2.2 with Airoboros 3.2.999https://huggingface.co/datasets/jondurbin/airoboros-3.2 Additionally, we incorporate the Aya dataset Singh et al. (2024) to enhance the multilingual capabilities of our base model. We filter out the instructions where language is not included in our pre-training language list. To bolster the model’s proficiency in Chinese, we include the COIG-CQIA Bai et al. (2024), ruozhiba-gpt4101010https://huggingface.co/datasets/hfl/ruozhiba_gpt4, and in-house Chinese multidisciplinary instruction data as supplementary datasets. To enhance math and coding abilities, we use the dart-math-hard Tong et al. (2024) and Magicoder-Evol-Instruct  [^11] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110KLuo et al. (2023) datasets.
@@
We leave the settings of post-training in Appendix A.
@@ 5 Experiments
We conducted extensive experiments to evaluate the capabilities of FuxiTranyu under the multilingual setting, specifically from the base model to the instruction-tuned model. We selected several models as benchmarks to compare our models with both English-centric and multilingual models. For English-centric models, we compared FuxiTranyu against Llama-2 (Llama-2-7B, Llama-2-chat-7B) (Touvron et al., 2023b) and Mistral (Mistral-7B-v0.1, Mistral-7B-instruct-v0.1) (Jiang et al., 2023a). For multilingual models, we compared FuxiTranyu with BLOOM (BLOOM-7B1, BLOOMZ-7B1) (Scao et al., 2022a; Muennighoff et al., 2022), PolyLM (PolyLM-13B, PolyLM-MultiAlpaca-13B) (Wei et al., 2023), and LLaMAX2 (LLaMAX2-7B, LLaMAX2-7B-Alpaca) (Lu et al., 2024). [^12] LLaMAX series models are continual pre-trained on the Llama-2 model to support beyond 100 languages. We used the LM Evaluation Harness framework (Gao et al., 2023) for all evaluation experiments.
@@
We evaluated the performance towards generative tasks, especially in translation and summarization tasks. For translation task, we employed WMT14 in en-fr translation direction (Bojar et al., 2014), WMT16 in en-de and en-ro translation directions (Bojar et al., 2016) and IWSLT 2017 (Cettolo et al., 2017) in en-ar translation direction for measuring the translation performance in our models and benchmark models. For summarization task, we used XL-Sum (Hasan et al., 2021) dataset. We selected 15 languages for the evaluation (ar, en, es, fr, gu, hi, id, mr, pt, ru, sr, ta, uk, vi, zh).
@@ 5.1 Base Model Evaluation
First, we report experiment results of our base models vs. baseline models. We focus on evaluating the capabilities of LLMs towards discriminative tasks. Evaluation results are shown in Table 5. Our model achieves the best performance on the XCOPA task. For other tasks, our model is significantly better than multilingual models like BLOOM-7B and PolyLM-13B. When compared to LLaMAX-7B, the evaluation results of our model are almost comparable, with no significant difference from the evaluation results of LLaMAX-7B. But compared with english-centric models, our model still worse than Llama-2-7B and Mistral-7B-v0.1 due to the limited training data used for English.
@@ 5.2 Instruction-Tuned Model Evaluation
We further compared our instruction-tuned model with other instruction-tuned models. We evaluated these models on both discriminative and generative tasks. Results are shown in Table 6. On discriminative tasks, our models achieve the best result on XCOPA. For m-Hellaswag, XWinograd, and XStoryCloze, our models outperforms the English-centric models, but slightly underperforms the multilingual models compared with LLaMAX2-7B. Our models still underperforms in m-ARC and m-MMLU tasks due to the limited training data used.
In generative tasks, our model excells on the summarization task, outperforming all baseline models. For the translation task, our model outperforms the English-centric models, but slightly underperforms the multilingual model like LLaMAX2-Alpaca-7B.
More details of our evaluations are discussed in Appendix B, where we report the results for each language tested.
@@ 6 Analysis and Interpretability
We further conducted an interpretability analysis of FuxiTranyu to provide a deep understanding of the underlying mechanisms driving its multilingual capabilities. To ensure a comprehensive analysis and consistency with prior research, we investigated our models from both the neuron Wu et al. (2023); Shi et al. (2024); Leng and Xiong (2024); Zhang et al. (2024); Tang et al. (2024); Liu et al. (2024); Kojima et al. (2024) and representation Conneau et al. (2020); Tiyajamorn et al. (2021); Chang et al. (2022); Rajaee and Pilehvar (2022); Xu et al. (2023); Dong et al. (2024); Xie et al. (2024) perspectives. Specifically, our neuron analysis explores the importance of different neurons to multilingual abilities of the model, while the representation analysis examines the characteristics of multilingual representations learned by the model. Here, we first introduce the details and results of our neuron analysis, while the representation analysis is discussed in Section 6.2.
@@ 6.1.1 Analysis Setup
6.1.2 Results
We analyzed the varying importance of different layers across diverse language inputs, as shown in Figure 4 (Appendix C). Our findings indicate that, universally, shallow layers exhibit low significance while deep layers demonstrate great importance. Notably, languages such as bn and ta exhibit a notably diminished importance in deep layers compared to others, aligning with our evaluation results where these languages perform poorly. This discrepancy may stem from their relatively limited representation learning in the pre-training data.
We then analyzed the significance of various components across different language inputs, depicted in Figure 5 (Appendix C), with 8 components per layer. Our findings mirror previous conclusions: components in shallow layers exhibit low importance, whereas those in deep layers show high significance. Moreover, a more detailed observation reveals that MLP components hold greater importance in shallow layers, whereas attention components are more critical in deep layers.
6.2 Representation Analysis
6.2.1 Analysis Setup
6.2.2 Results
Furthermore, we calculate the average similarity for each layer of the two models, as shown in Figure 6 (Appendix C). For our model, it can be observed that there is a significant increase in similarity from the embedding layer to layer 0, reaching a very high level. As the depth of the model increases, the similarity continues to rise, indicating that the model learns richer multilingual alignment information in these layers. Subsequently, there is a sharp decrease in similarity from layer 28 to layer 29, suggesting that language-specific multilingual representations in the final layer are learned to predict the diverse multilingual vocabulary. For BLOOM-7B1, the trend of similarity changes across layers is similar, initially increasing and then decreasing, but the changes are more gradual in magnitude.
@@ 7 Conclusion
In this paper, we have presented the FuxiTranyu models to address the need for open-source multilingual LLMs. Along with the base model, FuxiTranyu-8B, we also present the fine-tuned models on multilingual supervised fine-tuning dataset and preference dataset, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO. Evaluations on multilingual benchmarks show FuxiTranyu models outperform previous multilingual and monolingual LLMs. Furthermore, interpretability analyses underscore the efficacy of the multilingual capabilities embedded in FuxiTranyu.
@@ Acknowledgements
The present research was supported by the National Key Research and Development Program of China (Grant No. 2023YFE0116400). The computing resources used in this project are supported by the Scientific Computing Center of CIC, Tianjin University.
Appendix A Post-Training Details
@@
During the instruction tuning phase, we executed the fine-tuning process on 5 A100 80GB GPUs, leveraging the TRL framework for instruction fine-tuning and DPO training. Throughout both stages, we employed the ChatML format [^13] https://github.com/openai/openai-python/blob/release-v0.28.0/chatml.md for the chat template, and designated <PAD> as the pad token. We used AdamW Loshchilov and Hutter (2017) optimizer, complemented by a cosine learning rate scheduler. The maximum sequence length was set to 4096 for both stages.
@@
In the SFT stage, we configured the maximum learning rate to 2e-5, with a warmup phase spanning 10% of the total steps. The global batch size was set to 320, and the model was trained for 2 epochs. To optimize memory usage, we enabled Flash-Attention V2 Dao (2024), ZeRO stage 2 Rajbhandari et al. (2020), and gradient checkpointing. Additionally, we employed NEFTune Jain et al. (2023), which introduces noise to embedding weights to enhance the final performance of our instruction-tuned model.
@@
In the subsequent DPO training stage, we adhered to the latest hyper-parameters specified for reproducing the results of Zephyr, as provided by the alignment-handbook.141414alignment_handbook2023 The beta value for DPO was set to 0.01, and the training took 1 epoch on UltraFeedback. The maximum learning rate was set to 5e-7, with a warmup phase covering 10% of the total training steps. Similar to the SFT stage, the global batch size was maintained at 320, and we activated Flash-Attention V2 and gradient checkpointing to optimize memory usage. To accommodate the policy and reference model within memory constraints, we utilized ZeRO stage 3 for the policy model and omitted ZeRO for the reference model.
Appendix B Detailed Evaluation Results
@@
We provide detailed evaluation results for each language in this section. First, we present the results for all 15 tested languages on the multilingual ARC in Table 7, comparing base models and instruction-tuned models. The results show that our models perform better in 1 of the 15 tested languages for the ARC task. We speculate that our models still underforms on this task due to the relatively small amount of training data used.
@@
Instuction-tuned Model
@@
Instuction-tuned Model
@@
Next, we present the results for all 15 tested languages on multilingual HellaSwag in Table 8, comparing base models and instruction-tuned models. Despite our FuxiTranyu-8B model being trained on only about 600B tokens, it achieves remarkable performance. The SFT and RL-trained models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, also deliver promising results across all languages, even competing with powerful monolingual LLMs like Llama-2-7B and Mistral-7B-v0.1, with English language as exception.
@@
Instuction-tuned Model
Llama-2-Chat-7B 31.4 28.3 50.7 78.6 58.1 57.0 39.0 44.5
Mistral-7B-Instruct-v0.1 31.2 28.7 52.2 70.1 58.1 57.6 39.8 38.1
@@
Instuction-tuned Model
@@
We report results on multilingual MMLU in Table 9. Our models still underperforms baseline models for all languages. It is in line with the number of training tokens utilized in the pre-training process.
@@
Instuction-tuned Model
@@
Instuction-tuned Model
@@
Results on XWinograd are depicted in Table 10. Our FuxiTranyu SFT and DPO models achieve better results in Portuguese and Chinese. Although our models underperforms in English, French, Russian, and Japanese compared to Llama-2-7B, they outperforms previous multilingual LLMs like BLOOM-7B1 and PolyLM-13B across all languages.
@@
Instuction-tuned Model
@@
Results on XCOPA and XStoryCloze are shown in Table 11 and Table 12. For XCOPA, our base models achieve better results in sw, ta, tr, and vi. When compared to instruction-tuned models, our models achieve better results in more languages, specifically in it, id, ta, th, tr, vi, and zh. On the XStoryCloze task, our base models achieve better results in three languages: ar, my, and ru. However, for instruction-tuned models, our models outperforms other baseline models only in my.
@@
Instuction-tuned Model
@@
Instuction-tuned Model
@@
We present our evaluation results for generative tasks in Table 13 and Table 14. On the XL-Sum task, our models significantly outperform all baseline models across all evaluated languages, demonstrating the potential of our models for summarization task, particularly in a multilingual context. For the translation tasks in WMT14, WMT16, and IWSLT2017, our models excell in the en-ro, en-de, and en-fr translation directions. However, they still lag behind other baseline models in the ro-en, de-en, fr-en, ar-en, and en-ar translation directions. This indicates that our models perform significantly better for out-of-English translation directions. Although our models underperform in the en-ar direction compared to LLaMAX-2-Alpaca, they still achieve notably better results than other models.
@@
Table 13: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on XL-Sum (0-shot).
@@
Table 14: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, PolyLM-13B, and LLaMAX2-7B models on WMT14, WMT16, and IWSLT2017 (0-shot).
Appendix C Detailed Analysis Results

### v3

@@ 3.2 Model Architecture
  • Normalization. Given the significance of pre-training stability in training large LMs with a substantial number of tokens, we implement pre-normalization due to its superior stability compared to post-normalization Xiong et al. (2020). Furthermore, we incorporate the widely used RMSNorm Zhang and Sennrich (2019); Jiang et al. (2023b) to enhance training efficiency.
@@ 4 Post-training
To develop a model capable of following instructions and engaging in conversational interactions with humans, we have adopted the instruction fine-tuning and reinforcement learning (RL) approach outlined in Ouyang et al. (2022).
During the instruction fine-tuning phase, we curate a diverse and high-quality open-source instruction dataset. Given the abundance of instruction-following datasets that have demonstrated exceptional alignment results with various models, manually selecting and fine-tuning the mixture rates for each dataset becomes a challenging task. Consequently, we opt to designate a primary dataset and supplement it with additional datasets. In this context, we select the OpenHermes 2.5 data collection Teknium (2023) as our base dataset, which is composed of multiple datasets covering a wide range of instructions and yielding excellent results when fine-tuned with Mistral-7B-v0.1. We make modifications to the original OpenHermes 2.5 dataset by replacing Airoboros 2.2 with Airoboros 3.2.888https://huggingface.co/datasets/jondurbin/airoboros-3.2 Additionally, we incorporate the Aya dataset Singh et al. (2024) to enhance the multilingual capabilities of our base model. We filter out the instructions where language is not included in our pre-training language list. To bolster the model’s proficiency in Chinese, we include the COIG-CQIA Bai et al. (2024), ruozhiba-gpt4, [^9] https://huggingface.co/datasets/hfl/ruozhiba_gpt4 and in-house Chinese multidisciplinary instruction data as supplementary datasets. To enhance math and coding abilities, we use the dart-math-hard Tong et al. (2024) and Magicoder-Evol-Instruct  [^10] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110KLuo et al. (2023) datasets. The involved languages in the supervised fine-tuning stage can be found in Section C.2.
@@
We detail the settings of post-training in Section C.2.
@@ 5 Experiments
We conducted extensive experiments to evaluate the capabilities of FuxiTranyu (both the base model and instruction-tuned models) under the multilingual setting. We compared FuxiTranyu against strong baselines, including both English-centric and multilingual models. For English-centric models, we used Llama-2 (Llama-2-7B, Llama-2-chat-7B) (Touvron et al., 2023b) and Mistral (Mistral-7B-v0.1, Mistral-7B-instruct-v0.1) (Jiang et al., 2023a) as baseline. For multilingual models, we compared FuxiTranyu with BLOOM (BLOOM-7B1, BLOOMZ-7B1) (Scao et al., 2022a; Muennighoff et al., 2022), PolyLM (PolyLM-13B, PolyLM-MultiAlpaca-13B) (Wei et al., 2023), and LLaMAX2 (LLaMAX2-7B, LLaMAX2-7B-Alpaca) (Lu et al., 2024). [^11] LLaMAX series models are continual pre-trained on the Llama-2 model to support beyond 100 languages. We used the LM Evaluation Harness framework (Gao et al., 2023) for all evaluation experiments.
@@
We evaluated the performance towards generative tasks, especially in translation and summarization tasks. For the translation task, we employed WMT14 in en-fr translation direction (Bojar et al., 2014), WMT16 in en-de and en-ro translation directions (Bojar et al., 2016) and IWSLT 2017 (Cettolo et al., 2017) in en-ar translation direction for measuring the translation performance in our models and benchmark models. For the summarization task, we used XL-Sum (Hasan et al., 2021) dataset. We selected 15 languages for the evaluation (ar, en, es, fr, gu, hi, id, mr, pt, ru, sr, ta, uk, vi, zh).
@@ 5.1 Base Model Evaluation
First, we report the experiment results of our base models vs. baseline models. We focus on evaluating the capabilities of LLMs towards discriminative tasks. Evaluation results are shown in Table 2. Our model achieves the best performance on the XCOPA and XStoryCloze tasks. For other tasks, our model is significantly better than multilingual models like BLOOM-7B and PolyLM-13B. When compared to LLaMAX2-7B, the evaluation results of our model are almost comparable, with no significant difference from the evaluation results of LLaMAX2-7B. But compared with English-centric models, our model is still worse than Llama-2-7B and Mistral-7B-v0.1 due to the limited training data used for English.
@@ 5.2 Instruction-Tuned Model Evaluation
We further compared our instruction-tuned models with other instruction-tuned models. We evaluated these models on both discriminative and generative tasks. Results are shown in Table 3. On discriminative tasks, our models achieve the best result on XCOPA. For m-Hellaswag, XWinograd, and XStoryCloze, our models outperform the English-centric models but slightly underperform the multilingual models compared with LLaMAX2-7B. Our models still underperform in m-ARC and m-MMLU tasks due to the limited training data used.
In generative tasks, our models excel on the summarization task, outperforming all baseline models. For the translation task, our models outperform the English-centric models but slightly underperform the multilingual model like LLaMAX2-Alpaca-7B.
More details of our evaluations are discussed in Appendix D, where we report the results for each language tested.
@@ 6 Analysis and Interpretability
We further conducted an interpretability analysis of FuxiTranyu to provide a deep understanding of the underlying mechanisms driving its multilingual capabilities. To ensure a comprehensive analysis and consistency with prior research, we investigated our models from both the neuron Wu et al. (2023); Shi et al. (2024); Leng and Xiong (2024); Zhang et al. (2024); Tang et al. (2024); Liu et al. (2024); Kojima et al. (2024) and representation Conneau et al. (2020); Tiyajamorn et al. (2021); Chang et al. (2022); Rajaee and Pilehvar (2022); Xu et al. (2023); Dong et al. (2024); Xie et al. (2024) perspectives. Specifically, our neuron analysis explores the importance of different neurons to the multilingual abilities of the model, while the representation analysis examines the characteristics of multilingual representations learned by the model. Here, we first introduce the details and results of our neuron analysis, while the representation analysis is discussed in Section E.1.
@@ 6.2 Neuron Analysis Setup
6.3 Neuron Analysis Results
We analyzed the varying importance of different layers across diverse language inputs, as shown in Figure 4 (Section E.2). Our findings indicate that universally, shallow layers exhibit low significance while deep layers demonstrate great importance. Notably, languages such as bn and ta exhibit a notably diminished importance in deep layers compared to others, aligning with our evaluation results where these languages perform poorly. This discrepancy may stem from their relatively limited representation learning in the pre-training data.
We then analyzed the significance of various components across different language inputs, depicted in Figure 5 (Section E.2), with 8 components per layer. Our findings mirror previous conclusions: components in shallow layers exhibit low importance, whereas those in deep layers show high significance. Moreover, a more detailed observation reveals that MLP components hold greater importance in shallow layers, whereas attention components are more critical in deep layers.
@@ 7 Conclusion
In this paper, we have presented FuxiTranyu to address the need for open-source multilingual LLMs. Along with the base model, FuxiTranyu-8B, we also present instruction-tuned models fine-tuned on multilingual supervised fine-tuning and preference data, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO. Evaluations on multilingual benchmarks show FuxiTranyu outperforms previous multilingual and monolingual LLMs. Furthermore, interpretability analyses underscore the efficacy of the multilingual capabilities embedded in FuxiTranyu.
@@ Acknowledgements
The present research was supported by the National Key Research and Development Program of China (Grant No. 2023YFE0116400). The computing resources used in this project were supported by the Scientific Computing Center of the College of Intelligence and Computing, Tianjin University. We would like to thank the anonymous reviewers for their insightful comments.
Appendix A Supported Languages in FuxiTranyu
Appendix B Tokenization
Following the approach used in BLOOM, we also evaluate the performance of our tokenizer using the fertility metric. To assess its efficacy, we conduct a comparative analysis with the Llama-2 and BLOOM tokenizers. This evaluation involves computing fertility on the same set of documents across different languages. Results are presented in Figure 2, which indicate that the FuxiTranyu tokenizer is more efficient than the others in most languages. Based on our evaluations and interpretability analysis, we believe that the fertility of the tokenizer positively correlates with the model’s performance in specific languages. In the fertility test, we observe that Bengali (bn), Hindi (hi), and Tamil (ta) exhibit high fertility, indicating lower tokenization efficiency in these languages compared to others. Consequently, the performance and importance of neurons of these languages in our base model are also suboptimal. Further details are discussed in Section 6.3.
Appendix C Training Details
Pre-Training SFT DPO
# Params 8B Learning Rate 2e-5 Learning Rate 5e-7
Hidden Size 4,096 Warmup Ratio 10% Warmup Ratio 10%
Intermediate Size 16,384 Batch Size 512 Batch Size 512
Heads 32 Epochs 2 Epochs 1
Layers 30 NEFTune $\checkmark$ beta 0.01
FlashAttn V2 $\checkmark$ FlashAttn V2 $\checkmark$ FlashAttn V2 $\checkmark$
Training Tokens 606B # Instances 1M # Instances 61.1k
Position Embed 4,096
Vocab Size 250,752
Learning Rate 3e-4 $\rightarrow$ 1e-4
Batch Size 2M $\rightarrow$ 4M
Context Length 4,096
C.1 Pre-training Details
The training procedure for the FuxiTranyu model adheres to the standard autoregressive language model framework, utilizing the next-token prediction loss as detailed in (Brown et al., 2020). To enhance pre-training efficiency, we employ a document packing method similar to that described in (Raffel et al., 2020). This involves randomly shuffling documents, merging them, and then truncating them into multilingual chunks that adhere to a maximum context length of 4096 tokens during the pre-training phase.
To mitigate memory consumption and further improve training efficiency, we leverage ZeRO-2 Rajbhandari et al. (2020) and Flash-Attention V2 Dao (2024) technologies. For optimization, the standard AdamW optimizer Loshchilov and Hutter (2017) is utilized with hyper-parameters set to $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=10^{-8}$. We employ the cosine learning rate scheduler, starting with a maximum learning rate of 3e-4 and decaying to a minimum of 10% of the maximum rate. Notably, after encountering divergence issues when training approximately 241 billion tokens, we reduced the maximum learning rate to 1e-4 to match with the learning rate used in BLOOM, given the multilingual context of both models.
C.2 Post-Training Details
The instruction datasets collected do not cover all languages used during pre-training. For the current version of FuxiTranyu, we provide support for the following languages: Arabic, Bengali, Burmese, Chinese, Dutch, English, Filipino, Finnish, French, German, Greek, Hindi, Hungarian, Indonesian, Italian, Japanese, Korean, Kyrgyz, Malay, Persian, Polish, Portuguese, Russian, Spanish, Swedish, Tamil, Thai, Turkish, Ukrainian, Urdu, and Vietnamese.
@@ C.2 Post-Training Details
During the instruction tuning phase, we executed the fine-tuning process on 8 A800 80GB GPUs, leveraging the TRL framework for instruction fine-tuning and DPO training. Throughout both stages, we employed the ChatML format [^13] https://github.com/openai/openai-python/blob/release-v0.28.0/chatml.md for the chat template, and designated <PAD> as the pad token. We used AdamW Loshchilov and Hutter (2017) optimizer, complemented by a cosine learning rate scheduler. The maximum sequence length was set to 4096 for both stages.
@@
In the SFT stage, we configured the maximum learning rate to 2e-5, with a warmup phase spanning 10% of the total steps. The global batch size was set to 512, and the model was trained for 2 epochs. To optimize memory usage, we enabled Flash-Attention V2 Dao (2024), ZeRO stage 2 Rajbhandari et al. (2020), and gradient checkpointing. Additionally, we employed NEFTune Jain et al. (2023), which introduces noise to embedding weights to enhance the final performance of our instruction-tuned model.
@@
In the subsequent DPO training stage, we adhered to the latest hyper-parameters specified for reproducing the results of Zephyr, as provided by the alignment-handbook.141414alignment_handbook2023 The beta value for DPO was set to 0.01, and the training took 1 epoch on UltraFeedback. The maximum learning rate was set to 5e-7, with a warmup phase covering 10% of the total training steps. Similar to the SFT stage, the global batch size was maintained at 512, and we activated Flash-Attention V2 and gradient checkpointing to optimize memory usage. To accommodate the policy and reference model within memory constraints, we utilized ZeRO stage 3 for the policy model and omitted ZeRO for the reference model.
Appendix D Detailed Evaluation Results
@@
We provide detailed evaluation results for each language in this section. First, we present the results for all 15 tested languages on the multilingual ARC in Table 7, comparing base models and instruction-tuned models. In base models, the results show that our models perform better in 1 of the 15 tested languages for the ARC task. In instruction-tuned models, our models outperforms in ar and vi languages. We speculate that our models still underperformed on this task due to the relatively small amount of training data used.
@@
Instruction-tuned Model
@@
Instruction-tuned Model
@@
Next, we present the results for all 15 tested languages on multilingual HellaSwag in Table 8, comparing base models and instruction-tuned models. Despite our FuxiTranyu-8B model being trained on only about 600B tokens, it achieves remarkable performance. Comparing base models, our models outperforms other models in ar, bn, hu, and vi languages. Our model still lag behind compared with Mistral-7B-v0.1, but outperform other baseline models, except sk language. The SFT and RL-trained models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, also deliver promising results across all languages, even competing with powerful monolingual LLMs like Llama-2-7B and Mistral-7B-v0.1, with English and Spanish as an exception.
@@
[…]
Instruction-tuned Model
[…]
@@
Instruction-tuned Model
@@
We report results on multilingual MMLU in Table 9. Our models still underperform baseline models for all languages. It is in line with the number of training tokens utilized in the pre-training process.
@@
Instruction-tuned Model
@@
Instruction-tuned Model
@@
Results on XWinograd are depicted in Table 10. In base models, although our models still underperformed compared to Mistral-7B-v0.1, they outperforms previous multilingual LLMs like BLOOM-7B1 and PolyLM-13B across all languages. Notably, our FuxiTranyu SFT and DPO models achieve better results in Chinese.
@@
Instruction-tuned Model
@@
Results on XCOPA and XStoryCloze are shown in Table 11 and Table 12. For XCOPA, our base models achieve better results in sw, ta, tr, and vi. When compared to instruction-tuned models, our models achieve better results in more languages, specifically in it, ta, th, tr, and vi. On the XStoryCloze task, our base models achieve better results in three languages: ar, my, and ru. However, for instruction-tuned models, our models outperform other baseline models only in my.
@@
Instruction-tuned Model
@@
Instruction-tuned Model
@@
We present our evaluation results for generative tasks in Table 13 and Table 14. On the XL-Sum task, our models significantly outperform all baseline models across all evaluated languages, demonstrating the potential of our models on summarization task, particularly in a multilingual context. For the translation tasks in WMT14, WMT16, and IWSLT2017, our models excel in the en-ro and en-de translation directions. However, they still lag behind other baseline models in the ro-en, de-en, fr-en, ar-en, and en-ar translation directions. This indicates that our models perform significantly better for out-of-English translation directions. Although our models underperformed in the en-fr and en-ar directions compared to LLaMAX2-Alpaca, they still achieve notably better results than other models.
@@
Table 13: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, and LLaMAX2-7B models on XL-Sum (0-shot).
@@
FuxiTranyu-8B-DPO 31.98 59.64 32.27 60.19 27.05 51.5 6.5 29.41
Table 14: Performance of FuxiTranyu-8B models compared to Llama-2-7B, Mistral-7B-v0.1, BLOOM-7B1, and LLaMAX2-7B models on WMT14, WMT16, and IWSLT2017 (0-shot).
Appendix E Additional Analysis and Interpretability
E.1 Representation Analysis
E.1.1 Analysis Setup
E.1.2 Results
Furthermore, we calculate the average similarity for each layer of the two models, as shown in Figure 6 (Section E.2). For our model, it can be observed that there is a significant increase in similarity from the embedding layer to layer 0, reaching a very high level. As the depth of the model increases, the similarity continues to rise, indicating that the model learns richer multilingual alignment information in these layers. Subsequently, there is a sharp decrease in similarity from layer 28 to layer 29, suggesting that language-specific multilingual representations in the final layer are learned to predict the diverse multilingual vocabulary. For BLOOM-7B1, the trend of similarity changes across layers is similar, initially increasing and then decreasing, but the changes are more gradual in magnitude.
E.2 Detailed Analysis Results